OpenStudy (anonymous):

Convert the given function into a logarithmic function f(x)=8.71e^-0.55x

OpenStudy (anonymous):

I have this so far f(x)=8.71e^-0.55x y=8.71e 6-0.55x y/8.71 =e^-0.55x In y-In 8.71 = In e^-0.55x

OpenStudy (anonymous):

use rules of logs to combine the two terms on the right? and cancel the ln e?

OpenStudy (anonymous):

Can you show me how to do this?

OpenStudy (anonymous):

ln a - ln b = ln (a/b)

OpenStudy (anonymous):

ln e^a=a

sam (.sam.):

\[f(x)=8.71e^{-0.55x}\] let f(x)=y \[y=8.71e^{-0.55x} \\ \\ \ln(y)=\ln(8.71e^{-0.55x})\] Using ln(ab)=ln(a)+ln(b) \[\ln(y)=\ln(8.71)+\ln(e^{-0.55x})\] Also, \(ln(e^m)=m\), so \[\ln(y)=\ln(8.71)-0.55x\]